OpenStudy (lgbasallote):

Lgbariddle: I have a number in mind. You can guess, but you'll never find. If you try to take away one, Then take away seven more, It no longer is A number you're familiar of. But alas, if you add six, Then it becomes nine less Than it was before

mathslover (mathslover):

let the number be x x- 1 - 7 = let it be y x-8+6=y-10 y is a rational number... x-2 = y-10 x - y = -8

mathslover (mathslover):

M i right till this?

OpenStudy (lgbasallote):

i made a typo

mathslover (mathslover):

yuppi :)

OpenStudy (lgbasallote):

Alas you are wrong, You have to think more If this riddle, You shall solve

mathslover (mathslover):

ALAS ...

OpenStudy (lgbasallote):

Tick tock Think outside the box An engineer accepts truth, A mathematician sees lies

OpenStudy (lgbasallote):

you're researching my records...why...?

mathslover (mathslover):

What a researcher can do? Just re... searching

mathslover (mathslover):

:P

OpenStudy (lgbasallote):

i have more than 100 lgbariddles...you fail in your research

mathslover (mathslover):

let me re search :P ...

mathslover (mathslover):

well I am going too much off topic sorry

OpenStudy (anonymous):

u imagining imaginary number

OpenStudy (lgbasallote):

Alas, no This is real I have five fingers So do you

OpenStudy (lgbasallote):

He who hates Math Is he who knows not Math He who forgets Math Is he who knows Math